INAF LBC MODS queue observing, 2013 January 20-21

Observer: M.Faccini, F.Cusano
Telescope Operator: G.Bechetti
Support Astronomer: O. Khun
Instrument: LBC + MODS

Summary

At the beginning of the night thick clouds at the horizon, cirrus all over in the sky, moon 63% illumination, seeing 0.8. Starting with LBC. Problems with LBC, elongated stars in the blue channel. Trying to solve the problem. Mostly clear in the middle of the night. Switch to MODS.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Annibali 23B NGC 4449 dual grating 2h science target COMPLETED
Annibali 23B HD108177 dual grating 1s Lick standard COMPLETED
Annibali 23B Feige 66 dual grating   spec standard
Glli 18D Feige 66 red grating   spec standard
Glli 18D cxoj12 red grating 15m science target COMPLETED
Annibali 23A NGC 4449 Pn dual grating 2h science target

Details

START WITH LBC

01:39 slew to pointing field of CalchiNova 3A

01:47 dopfia first

01:58 launched M31_Focus.ob

02:01 elongated stars in both channel

02:05 changing z4 manually

02:06 dopfia X2

02:14 we try an exposure test with position angle zero, it did not work, stars still elongated in the blue channel

02:40 dopfia again *it's impossible to take science data in this condition *we decide to try to solve the problem due to the high ranking projects (Grazian 30) foreseen at the end of the night. Other MODS programs are in constraints at 7 UT currently we have only LBC targets *we are discussing the problem with the italians AO *we take images at different focus to understand if the features rotate

03:20 calling John Hill

03:44 dopfia

04:07 trying to correct manually (astigmatism)

04:17 dopfia and manual correction

05:44 slew to Clementini 21, Leo T

*05:50*After the test we reach a better quality for the images in a new field about 52deg
  • the results are not good for the Clementini field.
  • according with the situation and with the difficulties encountered (and also with the evaluation of the SA) we switch to MODS for the rest of the night.
  • the quality of the images is not useful for the Science prog. of this night *The technical data acquired during this test will be useful to solve the "elongation" problem for the next run.

06:17 SWITCH TO MODS

06:43 slew to science target of Annibali 23B

06:44 acqMods mods.1.4449runb.acq mods1r.20130121.0010/0012

06:59 execMODS mods.1.4449runb.obs mods1r.20130121.0013/0016, mods1b.20130121.0003/006

09:12 acqMods hd108177_lick_runb.acq mods1r.20130121.0017/0023

09:25 execMODS hd108177_lick_runb.obs mods1r.20130121.0024, mods1b.20130121.0007

09:37 acqMods feige56_std_runb.acq (standard for Annibali and Glli), mods1r.20130121.0025/0027

09:44 execMODS feige66_std_runb.acq (dual grating for Annibali and red grating for Gilli) mods1r.20130121.0028/0031, mods1b.20130121.0008 *error in the header of the file 0029-0031 the name of the star is feige110 while is feige66

09:58 slew to Gilli 18D, acqMods cxoj12.acq mods1r.20130121.0032/0034

10:11 execMODS cxoj12.obs mods1r.20130121.0035

10:31 slew to Annibali 23A, acqMods mods.4449runa.acq mods1r.20130121.0036/38

10:51 execMODS mods.4449runa.obs mods1r.20130121.0039/41, mods1b.20130121.0009/11

13:28 starting calibration

-- MarcoFaccini - 21 Jan 2013
Topic revision: r6 - 29 Jan 2013, RobertoSpeziali
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback